OpenStudy (anonymous):

What is the solution to 3.2x = 6.4? A. x=2 B. x = 3.2 C. x = 9.6 D. x = 20.48

OpenStudy (matt101):

Divide both sides by 3.2 to isolate x: 3.2x/3.2=6.4/3.2 x=2 Also posting this in the math section will get you a quicker answer next time :)